Derive an equation to calculate the income elasticity of demand at any point on the following function: Q = 40 - Y-3/2
1
Expert's answer
2020-05-25T09:52:44-0400

The income elasticity of demand at any point on the function Q = 40 - Y^(-3/2) is:

Ei = -b×Y/Q = 3/2×Y/Q.